T1=1,T1=2Tn-1+2
write the first five terms of the sequence

Answer by Edwin McCravy(20060)

write the first five terms of the sequence
T1 = 1, Tn = 2Tn-1 + 2


You already have T1, so you replace n by 2, 
then by 3, then by 4, then by 5.

Tn = 2Tn-1 + 2

Replacing n by 2:

T2 = 2T2-1+2 = 2T1+2 = 2(1)+2 = 2+2 = 4

Replacing n by 3:

T3 = 2T3-1+2 = 2T2+2 = 2(4)+2 = 8+2 = 10

Replacing n by 4:

T4 = 2T4-1+2 = 2T3+2 = 2(10)+2 = 20+2 = 22

Replacing n by 5:

T5 = 2T5-1+2 = 2T4+2 = 2(22)+2 = 44+2 = 46

and the sequence is

T1, T2, T3, T4, T5

or  1, 4, 10, 22, 46

Edwin